# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 9. Torsion-free modules over $h$-local domains
Now we are ready to apply our results to $\mbox{\rm Add}(M)$ for $M$ a finitely generated torsion-free module $M$ over an $h$-local domain. That is, we are going to apply the previous results to $\Lambda=\operatorname{End}_{R}(M)$. We need to keep in mind the following lemma.
@@
    For any non-zero two-sided ideal $I$ of $\Lambda$, there are canonical isomorphisms
    $R/(I\cap R)\longrightarrow(R/(I\cap R))_{\mathfrak{m}_{1}}\times\dots\times(R/(I\cap R))_{\mathfrak{m}_{t}}$
    and
    $\Lambda/I\longrightarrow(\Lambda/I)_{\mathfrak{m}_{1}}\times\dots\times(\Lambda/I)_{\mathfrak{m}_{t}},$
    where $\{\mathfrak{m}_{1},\dotsc,\mathfrak{m}_{t}\}$ are the unique ideals of $R$ containing $I\cap R$.
So let $M_{R}$ be a finitely generated torsion-free module of rank $k$ over an $h$-local domain $R$. Let $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ be the list of those maximal ideals of $R$ over which $M_{\mathfrak{m}}$ is not free. Assume further that $\Lambda$ is locally semiperfect, that is, that each $M_{\mathfrak{m}}$ is a direct sum of modules with local endomorphism rings. For $i=1,\dots,\ell$, let $M_{\mathfrak{m}_{i}}\cong\bigoplus_{j=1}^{t_{i}}M_{i,j}^{m_{i,j}}$, where $M_{i,1},\dots,M_{i,t_{i}}$ are pairwise non-isomorphic indecomposable modules and $m_{i,j}\in\mathbb{N}$. Note that $\Lambda_{\mathfrak{m}_{i}}\cong\bigoplus_{j=1}^{t_{i}}U_{i,j}^{m_{i,j}}$, where $U_{i,j}=\operatorname{Hom}_{R_{\mathfrak{m}_{i}}}(M_{\mathfrak{m}_{i}},M_{i,j})$, is an indecomposable decomposition of $\Lambda_{\mathfrak{m}_{i}}$. Let $r_{i,j}$ be the length of $(U_{i,j})_{Q}$. As a $Q$-module, $(U_{i,j})_{Q}=Q^{kr_{i,j}}$. On the other hand, we also have isomorphisms of $Q$-modules $(U_{i,j})_{Q}\cong\operatorname{Hom}_{Q}(M_{Q},(M_{i,j})_{Q})\cong(M_{i,j})_{Q}^{k}=Q^{k\operatorname{rank}(M_{i,j})}$. Therefore $r_{i,j}=\operatorname{rank}(M_{i,j})$.
@@
[…]
@@
    for every $\mathfrak{m}\in\operatorname{mSpec}R$, $M_{\mathfrak{m}}$ contains a direct summand isomorphic to $R_{\mathfrak{m}}$.
Proof.  The assumption $\operatorname{Tr}_{R}(M)=R$ is equivalent to $R\in\mbox{\rm add}(M)$. In particular, $\mbox{\rm Add}(M)$ contains $R^{(\omega)}\oplus M$, so $(i)$ follows from [AHP, Proposition 6.1]. Note that $\operatorname{Tr}_{R_{\mathfrak{m}}}(M_{\mathfrak{m}})=R_{\mathfrak{m}}$ for any maximal ideal $\mathfrak{m}\in\operatorname{mSpec}R$, hence $R_{\mathfrak{m}}\in\mbox{\rm add}(M_{\mathfrak{m}})$. Since $R_{\mathfrak{m}}$ is local, $R_{\mathfrak{m}}$ has to be isomorphic to a direct summand of $M_{\mathfrak{m}}$.
@@
Let $R$ be an $h$-local domain and let $M$ be a non-zero finitely generated torsion-free $R$-module with endomorphism ring $\Lambda=\operatorname{End}_{R}(M)$. Consider the following statements,
@@
    $\Lambda$ is locally semiperfect and if $X$ and $Y$ are indecomposable direct summands of $M_{\mathfrak{m}}$ and $N_{\mathfrak{n}}$ for some $\mathfrak{m}\neq\mathfrak{n}\in\operatorname{mSpec}R$, then $\operatorname{rank}(X)$ and $\operatorname{rank}(Y)$ are coprime.
@@
Let us show that $(i)\Rightarrow(ii)$. Assume that $(i)$ is true. Note that $r_{i,a}$ and $r_{j,b}$ are coprime whenever $i\neq j$, $a\in\{1,\dots,t_{i}\}$, $b\in\{1,\dots,t_{j}\}$. We want to check that $(ii)$ of Proposition 8.3 is true.
Consider $a_{1},\dots,a_{\ell}$, $a_{i}\in\{1,\dots,t_{i}\}$ and $A=\prod_{i=1}^{\ell}r_{i,a_{i}}$. Fix $1\leq j\leq\ell$. Since $r_{1,a_{1}},\dots,r_{\ell,a_{\ell}}$ are pairwise coprime, there exists $t\in\{0,\dots,A-1\}$ such that $t\equiv 0\ ({\rm mod}\ r_{i,a_{i}})$ if $i\neq j$ and $t\equiv r_{j,b}\ ({\rm mod}\ r_{j,a_{j}})$. If $t<r_{j,b}$ we can increase $t$ by a suitable multiple of $A$, so we may assume $t\geq r_{j,b}$. Set $x_{i}:=t/r_{i,a_{i}}$ if $i\neq j$ and $x_{j}:=(t-r_{j,b})/r_{j,a_{j}}$. Then $x_{1},\dots,x_{\ell}$ are non-negative integers and $t=x_{i}r_{i,a_{i}}$ if $i\neq j$ and $t=x_{j}r_{j,a_{j}}+r_{j,b}$. In other words, the condition $(ii)$ of Proposition 8.3 is true. Then any projective $\Lambda$-module is a direct sum of finitely generated modules. $(ii)$ then follows from Theorem 4.11.
Now, assume that $(ii)$ is true, and suppose that $\operatorname{Tr}_{R}(M)=R$, and $R$ has Krull dimension $1$. Then, by Lemma 9.2, $\Lambda$ is locally semiperfect. Keeping the notation of Proposition 8.3, we have to prove that $r_{i,a}$ and $r_{j,b}$ are coprime whenever $i\neq j$, $a\in\{1,\dots,t_{i}\}$, $b\in\{1,\dots,t_{j}\}$. Note that it is trivial if either $r_{i,a}=1$ or $r_{j,b}=1$. On the other hand, if $r_{i,a}>1$ then there exists $c\in\{1,\dots,t_{i}\}$ such that $r_{i,c}=1$. Then $r_{i,c}+xr_{i,a}=yr_{j,b}$ can be solvable only if $r_{i,a},r_{j,b}$ are coprime. Now $(i)$ follows from Proposition 8.3.

### v2

@@ 9. Torsion-free modules over $h$-local domains
In this section we will use extensively the notation introduced in § 4.3 and Proposition 4.11. We apply the previous results to $\mbox{\rm Add}(M)$ for $M$ a finitely generated torsion-free module $M$ over an $h$-local domain.
We need to keep in mind the following lemma.
@@
    For any non-zero two-sided ideal $I$ of $\Lambda$, $I\neq\Lambda$, there is a canonical isomorphism $\Lambda/I\longrightarrow(\Lambda/I)_{\mathfrak{m}_{1}}\times\dots\times(\Lambda/I)_{\mathfrak{m}_{t}}$, where $\{\mathfrak{m}_{1},\dotsc,\mathfrak{m}_{t}\}$ are the unique maximal ideals of $R$ containing $I\cap R$.
Let $M_{R}$ be a finitely generated torsion-free module of rank $k$ over an $h$-local domain $R$. Following Lemma 9.1, let $\mathfrak{m}_{1},\dots,\mathfrak{m}_{\ell}$ be the list of those maximal ideals of $R$ over which $M_{\mathfrak{m}}$ is not free. Assume further that $\Lambda$ is locally semiperfect. That is, each $M_{\mathfrak{m}}$ is a direct sum of modules with local endomorphism rings, cf. Example 6.6. For $i=1,\dots,\ell$, let $M_{\mathfrak{m}_{i}}\cong\bigoplus_{j=1}^{t_{i}}M_{i,j}^{m_{i,j}}$, where $M_{i,1},\dots,M_{i,t_{i}}$ are pairwise non-isomorphic indecomposable modules and $m_{i,j}\in\mathbb{N}$. By Proposition 4.11, $\Lambda_{\mathfrak{m}_{i}}\cong\bigoplus_{j=1}^{t_{i}}U_{i,j}^{m_{i,j}}$, where $U_{i,j}=\operatorname{Hom}_{R_{\mathfrak{m}_{i}}}(M_{\mathfrak{m}_{i}},M_{i,j})$ is an indecomposable decomposition of $\Lambda_{\mathfrak{m}_{i}}$. Let $r_{i,j}$ be the length of $(U_{i,j})_{Q}$. As a $Q$-module, $(U_{i,j})_{Q}=Q^{kr_{i,j}}$. On the other hand, by Lemma 5.2, $(U_{i,j})_{Q}\cong\operatorname{Hom}_{Q}(M_{Q},(M_{i,j})_{Q})\cong(M_{i,j})_{Q}^{k}=Q^{k\operatorname{rank}(M_{i,j})}$. Therefore, $r_{i,j}=\operatorname{rank}(M_{i,j})$.
@@
Let $R$ be an $h$-local domain of Krull dimension $1$, and let $M$ be a finitely generated torsion-free $R$-module which is a generator of $\mbox{\rm Mod-}R$, i.e., $\operatorname{Tr}_{R}(M)=R$. Assume that every module of $\mbox{\rm Add}(M)$ is a direct sum of finitely generated modules. Then
@@
    for every $\mathfrak{m}\in\operatorname{mSpec}(R)$, $M_{\mathfrak{m}}$ contains a direct summand isomorphic to $R_{\mathfrak{m}}$.
Proof. The assumption $\operatorname{Tr}_{R}(M)=R$ is equivalent to $R\in\mbox{\rm add}(M)$. In particular, $\mbox{\rm Add}(M)$ contains $R^{(\omega)}\oplus M$, so $(i)$ follows from [AHP, Proposition 6.1]. By Lemma 5.3, $\operatorname{Tr}_{R_{\mathfrak{m}}}(M_{\mathfrak{m}})=R_{\mathfrak{m}}$ for any maximal ideal $\mathfrak{m}\in\operatorname{mSpec}(R)$, hence $R_{\mathfrak{m}}\in\mbox{\rm add}(M_{\mathfrak{m}})$. Since $R_{\mathfrak{m}}$ is local, $R_{\mathfrak{m}}$ has to be isomorphic to a direct summand of $M_{\mathfrak{m}}$ (otherwise, $\operatorname{Tr}_{R_{\mathfrak{m}}}(M_{\mathfrak{m}})\subseteq\mathfrak{m}R_{\mathfrak{m}}$).
@@
Let $R$ be an $h$-local domain, and let $M$ be a non-zero finitely generated torsion-free $R$-module with endomorphism ring $\Lambda=\operatorname{End}_{R}(M)$. Consider the following statements,
@@
    $\Lambda$ is locally semiperfect and if $X$ and $Y$ are indecomposable direct summands of $M_{\mathfrak{m}}$ and $M_{\mathfrak{n}}$, respectively, for some $\mathfrak{m}\neq\mathfrak{n}\in\operatorname{mSpec}(R)$. Then $\operatorname{rank}(X)$ and $\operatorname{rank}(Y)$ are coprime.
@@
Let us show that $(i)\Rightarrow(ii)$. Assume that $(i)$ is true. We will use the notation introduced after Lemma 9.1.
Note that $r_{i,a}$ and $r_{j,b}$ are coprime whenever $i\neq j$, $a\in\{1,\dots,t_{i}\}$, $b\in\{1,\dots,t_{j}\}$. We want to check that $(ii)$ of Proposition 8.3 is true.
Consider $a_{1},\dots,a_{\ell}$, $a_{i}\in\{1,\dots,t_{i}\}$, and $A=\prod_{i=1}^{\ell}r_{i,a_{i}}$. Fix $1\leq j\leq\ell$. Since $r_{1,a_{1}},\dots,r_{\ell,a_{\ell}}$ are pairwise coprime, there exists $t\in\{0,\dots,A-1\}$ such that $t\equiv 0\ ({\rm mod}\ r_{i,a_{i}})$ if $i\neq j$ and $t\equiv r_{j,b}\ ({\rm mod}\ r_{j,a_{j}})$. If $t<r_{j,b}$ we can increase $t$ by a suitable multiple of $A$, so we may assume $t\geq r_{j,b}$. Set $x_{i}:=t/r_{i,a_{i}}$ if $i\neq j$ and $x_{j}:=(t-r_{j,b})/r_{j,a_{j}}$. Then $x_{1},\dots,x_{\ell}$ are non-negative integers and $t=x_{i}r_{i,a_{i}}$ if $i\neq j$ and $t=x_{j}r_{j,a_{j}}+r_{j,b}$. In other words, the condition $(ii)$ of Proposition 8.3 is true. Then any projective $\Lambda$-module is a direct sum of finitely generated modules. $(ii)$ then follows from Proposition 4.11.
Now, assume that $(ii)$ is true, and suppose that $\operatorname{Tr}_{R}(M)=R$, and $R$ has Krull dimension $1$. Then, by Lemma 9.2, $\Lambda$ is locally semiperfect. Keeping the notation of Proposition 8.3, we have to prove that $r_{i,a}$ and $r_{j,b}$ are coprime whenever $i\neq j$, $a\in\{1,\dots,t_{i}\}$, and $b\in\{1,\dots,t_{j}\}$. Note that it is trivial if either $r_{i,a}=1$ or $r_{j,b}=1$. By Lemma 9.2, for any $\mathfrak{m}\in\operatorname{mSpec}(R)$, $R_{\mathfrak{m}}$ is a direct summand of $M_{\mathfrak{m}}$ (so $M_{\mathfrak{m}}$ has a direct summand of rank $1$, cf. the remarks before Lemma 9.2). Therefore, if $r_{i,a}>1$, then there exists $c\in\{1,\dots,t_{i}\}$ such that $r_{i,c}=1$. Then $r_{i,c}+xr_{i,a}=yr_{j,b}$ is solvable only if $r_{i,a},r_{j,b}$ are coprime. Now $(i)$ follows from Proposition 8.3.